# ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation

### v1

@@ Abstract
Reranking documents based on their relevance to a given query is a critical task in information retrieval. Traditional reranking methods often focus on improving the initial rankings but tend to lack transparency, failing to explain why one document is ranked higher than another. In this paper, we introduce ReasoningRank, a novel reranking approach that enhances clarity by generating two types of reasoning: explicit reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another. We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, more resource-efficient student models. While the student models may not outperform LLMs in speed, they significantly reduce the computational burden by requiring fewer resources, making them more suitable for large-scale or resource-constrained settings. These student models are trained to both generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT. Experiments demonstrate that ReasoningRank improves reranking accuracy and provides valuable insights into the decision-making process, offering a structured and interpretable solution for reranking tasks.  [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation
@@ 1 Introduction
The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al. (2023b); Sun et al. (2023). Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions. Recent studies have demonstrated that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al. (2023b); Liu et al. (2024); Yu et al. (2024c); Sun et al. (2023). By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
However, most existing reranking models do not distinguish between two key types of reasoning: explicit reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another. This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
Figure 1: Illustration of the two types of reasoning: explicit reasoning provides direct answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
In this paper, we propose ReasoningRank, a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings. Our approach is based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings. These explanations are then distilled into smaller, more computationally efficient student model that perform the reranking tasks. Our contributions are threefold:
  • We introduce a reranking approach that separates explicit and comparative reasoning, enhancing clarity of document rankings.
  • We demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
  • We provide an open-source implementation to facilitate further research and applications in information retrieval.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 presents the proposed ReasoningRank framework. Section 4 not only presents the quantitative performance metrics but also provides a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability. Finally, Section 6 concludes the paper and discusses future work.
@@ 2 Related Work
Figure 2: Overview of explicit and comparison reasoning prompts for document ranking. Explicit reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
@@
Figure 3: Overview of the ReasoningRank framework. The teacher model generates explicit and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
@@
Early ranking systems primarily focused on pointwise methods, where each document’s relevance was judged in isolation Liu et al. (2009). However, pairwise approaches like duoBERT Pradeep et al. (2021) were introduced to compare pairs of documents, determining which document was more relevant. While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list. To address this, listwise models such as RankGPT Sun et al. (2023) and its distilled versions like RankVicuna Pradeep et al. (2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy. However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
@@
The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al. (2009). Pointwise approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs. On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking. Despite these advances, most early methods lacked clarity that modern applications demand.
@@
Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions. Models such as Explain-then-Rank Zhuang et al. (2023) have focused on generating explanations that describe how a document matches a query. However, these approaches often only address explicit reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher). Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations. This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al. (2024); Yu et al. (2024b, a).
@@
Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al. (2015) to further reduce the computational demands of large reranking models. Knowledge distillation has been particularly effective in reranking tasks, as demonstrated by models like RankVicuna Pradeep et al. (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance. Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al. (2023); Dai et al. (2022); Xu et al. (2024b, a).
@@
By integrating these advancements, our ReasoningRank framework improves upon existing methods by offering a dual reasoning strategy—explicit and comparative reasoning—within a computationally efficient model. Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
@@ 3 Method
The overview of the architecture of the ReasoningRank model is shown in Figure 3. We propose ReasoningRank, a framework that integrates explicit and comparative reasoning within a unified reranking approach.
@@ 3.1 Teacher Model with Dual Reasoning
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both explicit and comparative aspects, and the examples can be seen in Figure 2: Explicit Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
@@ 3.2 Student Model with Knowledge Distillation
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we designed the student model to replicate the teacher’s performance while being more efficient for real-world deployment. The innovation lies in using knowledge distillation: the student learns from the teacher’s detailed reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
Architecture and Training:
@@
3.2.1 Training Objective and Loss Functions
To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model. First, using pairwise loss to optimize the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher. The pairwise loss is defined as:
@@ 3.2.1 Training Objective and Loss Functions
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence to measure the difference between the predicted ranking distribution and the target ranking distribution:
@@
Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. Based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
@@
where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground truth reasoning label.
3.3 Final Training Objective
@@
4 Experiments and Results
In this section, we evaluate the performance of our proposed ReasoningRank framework across various information retrieval tasks. We introduce the datasets and evaluation metrics used in our experiments (§4.1). We then detail the experimental setup and baseline methods for comparison (§4.3). Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in (§4.4). We performed paired t-tests across all queries in each dataset to assess the statistical significance of the improvements. The annotations in Tables 1 and 2 indicate the significance levels.
@@ 4.1 Datasets and Evaluation Metrics
BLEU, ROUGE-L, and Coherence Score: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
4.2 Teacher Models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4: A state-of-the-art LLM known for its strong reasoning and language understanding capabilities. ClaudeThe : An advanced LLM developed by Anthropic, offering competitive performance in reasoning tasks, we used the version claude-3-5-sonnet-20240620. GeminiTeam et al. (2023): A commercial LLM used for comparison purposes.We used the version gemini-1.5-flash.
4.2.1 Overall Performance on Benchmark Datasets
Table 1 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.69% on DL19 and 85.27% on TREC-COVID when using both comparison and explicit reasoning prompts, showcasing its robust handling of complex queries.
Adding different reasoning prompts to the LLMs enhances their performance across various datasets. For instance, GPT-4 with comparison and explicit reasoning outperforms its baseline (without reasoning) on DL19 by 2.10%. Similar trends are observed with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial.
To confirm the statistical significance of the improvements, we conducted paired t-tests comparing models with reasoning prompts to their respective baselines. The results, indicated by and in Table 1, show that the enhancements are significant, with p-values less than 0.05.
Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
BM25 50.58 47.96 59.47 44.22 39.52 30.75 40.73 31.83
Rank GPT-4 Models
Rank GPT-4 (without reasoning) 75.59 70.56 85.51 38.47 57.55 36.62 57.55 47.12
+ Explicit reasoning 75.32 69.93 83.27 38.94 55.87 36.08 52.89 46.98
+ Comparison reasoning 76.25 71.07 84.62 38.23 56.49 36.52 56.53 48.15
+ Comparison & Explicit 77.69 73.24 85.27 38.29 58.43 36.33 58.63 49.52
Claude Models
Claude (without reasoning) 71.96 68.81 82.77 36.12 54.35 36.36 50.78 42.51
+ Explicit reasoning 70.59 68.23 82.22 35.92 55.43 35.59 52.38 42.51
+ Comparison reasoning 71.34 69.75 83.31 36.43 55.11 36.71 51.53 43.23
+ Comparison & Explicit 72.13 69.95 84.03 37.33 55.52 36.36 52.67 44.91
Gemini Models
Gemini (without reasoning) 71.10 68.45 82.95 35.62 53.65 35.49 50.12 42.92
+ Explicit reasoning 69.08 67.02 81.55 34.83 52.87 36.92 51.31 43.92
+ Comparison reasoning 70.17 68.24 82.41 35.91 53.35 35.42 51.12 42.85
+ Comparison & Explicit 71.35 68.57 83.47 36.97 53.80 36.12 51.78 43.67
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
Student Models
Basic w/o reasoning 73.23 70.80 79.94 35.10 51.90 34.60 52.82 46.31
With explicit reasoning 74.54 70.11 79.37 34.25 52.81 35.45 50.75 46.95
With comparison reasoning 74.12 72.25 80.10 34.00 53.10 36.63 52.11 47.21
Comparison & Explicit 75.39 72.36 84.59 36.22 53.81 36.37 53.54 47.85
Baseline Models
DeBERTa Sun et al. (2023) 68.53 64.18 73.42 32.13 50.24 33.67 49.23 45.39
MonoT5 Nogueira et al. (2020) 74.54 70.44 79.95 34.07 46.03 35.21
RankVicuna Pradeep et al. (2023a) 71.12 68.66 67.14 48.73 38.53 55.67 35.29
RankZephyr Pradeep et al. (2023b) 72.23 70.53 85.12 36.46 53.33 38.92 60.72 35.52
APEER Jin et al. (2024) 74.57 72.25 83.86 35.25 52.11 33.41 56.03 46.13
Table 2: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with "–" indicate that results are not available for that dataset.
4.3 Student Models
Experimental Setup

### v2

# Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation

@@ Abstract
Reranking documents based on their relevance to a given query is a critical task in information retrieval. Traditional reranking methods often lack transparency and rely on proprietary models, hindering reproducibility and interpretability. We propose Reason-to-Rank (R2R), a novel open-source reranking approach that enhances transparency by generating two types of reasoning: direct relevance reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another. We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, openly available student models. Our student models are trained to generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT. Experiments demonstrate that R2R not only improves reranking accuracy but also provides valuable insights into the decision-making process. By offering a structured and interpretable solution with openly accessible resources, R2R aims to bridge the gap between effectiveness and transparency in information retrieval, fostering reproducibility and further research in the field. [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
@@ 1 Introduction
The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al. (2023b); Sun et al. (2023). Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions. Recent studies have revealed that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al. (2023b); Liu et al. (2024); Yu et al. (2024c); Sun et al. (2023). By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
However, most existing reranking models do not distinguish between two key types of reasoning: direct relevance reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another. This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
Figure 1: Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
In this paper, we propose Reason-to-Rank (R2R), a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings. Our approach based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings. Distilling these explanations into smaller, more computationally efficient student model that perform the reranking tasks. Our contributions are threefold:
  • Introduce a reranking approach that separates direct relevance and comparative reasoning, enhancing the clarity of document rankings.
  • Demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
  • Provide an open-source implementation to facilitate further research and applications in information retrieval.
The remainder of this paper is organize as follows. §2 reviews related work. §3 presents the Reason-to-Rank (R2R) framework. Sections 4 not only present the quantitative performance metrics but also provide a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability. Finally, §6 concludes the paper and discusses future work.
@@ 2 Related Work
Figure 2: Overview of direct relevance and comparison reasoning prompts for document ranking. Pairwise reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
@@
Figure 3: Overview of the Reason-to-Rank framework. The teacher model generates direct relevance and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
@@
Early ranking systems primarily focused on direct relevance methods, where each document’s relevance was judged in isolation Liu et al. (2009). However, pairwise approaches like duoBERT Pradeep et al. (2021) were introduced to compare pairs of documents, determining which document was more relevant. While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list. To address this, listwise models such as RankGPT Sun et al. (2023) and its distilled versions like RankVicuna Pradeep et al. (2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy. However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
@@
The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al. (2009). Direct relevance approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs. On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking. Despite these advances, most early methods lacked clarity that modern applications demand.
@@
Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions. Models such as Explain-then-Rank Zhuang et al. (2023) have focused on generating explanations that describe how a document matches a query. However, these approaches often only address direct relevance reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher). Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations. This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al. (2024); Yu et al. (2024b, a).
@@
Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al. (2015) to further reduce the computational demands of large reranking models. Knowledge distillation has been particularly effective in reranking tasks, as shown by models like RankVicuna Pradeep et al. (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance. Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al. (2023); Dai et al. (2022); Xu et al. (2024b, a).
@@
By integrating these advances, our R2R framework improves existing methods by offering a dual reasoning strategy - direct relevance and comparative reasoning - within a computationally efficient model. Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
@@ 3 Method
The overview of the architecture of the R2R model is shown in Figure 3. R2R, a framework that integrates direct relevance and comparative reasoning within a unified reranking approach.
@@ 3.1 Teacher Model with Dual Reasoning
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both direct relevance and comparative aspects, and the examples can be seen in Figure 2: Direct relevance Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
@@ 3.2 Student Model with Knowledge Distillation
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we design the student model to replicate the teacher’s performance while being more efficient for real-world deployment. The innovation lies in using knowledge distillation: the student learns from the teacher’s reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
3.2.1 Model Overview
@@
3.2.2 Training Objective
To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model. First, the pairwise loss optimizes the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher. The pairwise loss is defined as:
@@ 3.2.2 Training Objective
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence Hershey and Olsen (2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
@@
Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. It is based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
@@
where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground-truth reasoning label.
@@
Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
Baseline Models
DeBERTa Sun et al. (2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
MonoT5 Nogueira et al. (2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
RankVicuna Pradeep et al. (2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
RankZephyr Pradeep et al. (2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
APEER Jin et al. (2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
Our Models (student)
w/o reasoning (Pairwise Loss) 73.2 70.8 50.1 79.9 35.1 51.9 34.6 52.8 46.3
w/o reasoning (Listwise Loss) 73.8 71.0 50.5 80.2 35.4 52.3 35.0 53.0 46.6
w/ direct relevance reasoning 74.5 70.1 49.8 79.4 34.3 52.8 35.5 50.8 47.0
w/ comparison reasoning 74.1 72.3 50.9 80.1 34.0 53.1 36.6 52.1 47.2
w/ direct relevance & comparison 75.4 72.4 52.4 84.6 36.2 53.8 36.4 53.5 47.9
Table 1: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. ’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions. ’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations. ’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with ’–’ indicate that results are not available for that dataset.
4 Experiments
In this section, we evaluate the performance of our Reason-to-Rank (R2R) framework across various information retrieval tasks. We introduce the datasets and evaluation metrics used in our experiments §4.1. We then detail the experimental setup and baseline methods for comparison §4.2. Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in §4.2.4 and §4.2.1.
@@ 4.1 Datasets and Evaluation Metrics
BLEU, and ROUGE-L: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
4.2 Models
4.2.1 Teacher models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 Achiam et al. (2023), Claude The (claude-3-5-sonnet-20240620) and Gemini Team et al. (2023) (gemini-1.5-flash). These advanced LLMs known for their strong reasoning and language-understanding capabilities.
Teacher Model Performance
Table 5 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.6% on DL19 and 85.2% on TREC-COVID when using both comparison and direct relevance reasoning prompts, showcasing its robust handling of complex queries.
Adding different reasoning prompts to the LLMs enhances their performance across various datasets. For instance, GPT-4 with comparison and direct relevance reasoning outperforms its baseline (without reasoning) on DL19 by 2.1%. Similar trends are observesd with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial. To confirm the statistical significance of the improvements, we conduct paired t-tests comparing models with reasoning prompts to their respective baselines. Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
4.2.2 Student models